# Koopman for stochastic dynamics: error bounds for kernel extended dynamic mode decomposition

### v1

@@ 1. Introduction
Nowadays, the Koopman operator is an established mathematical and computational tool for modeling, analysis, and control of dynamical systems, see, e.g., the collection [MaurSusu20] or the recent overview papers [brunton:budisic:kaiser:kutz:2022, strasser2026overview]. The underlying idea is to consider nonlinear dynamics through the lens of observable functions using the linear, but infinite-dimensional Koopman operator, see the original work by B.O. Koopman [Koop31, KoopNeum32] and the seminal paper [Igor2005] by I. Mezić. This enabled, among others, to leverage modern techniques from spectral analysis to investigate the complex dynamical behaviour of the original nonlinear system [rowley2009spectral].
@@
Recently, data-driven approaches like dynamic mode decomposition (EDMD; [WillKevr15, Colb23]) were proposed to approximate the Koopman operator, see, e.g., the recent overview [Colb24], which has lead to a plethora of successful applications, including fluid flows [bagheri:2013, mezic:2013], climate forecasting [hogg:fonoberova:mezic:2020, azencot:erichson:lin:mahoney:2020], molecular dynamics [nuske:klus:2023], power systems technology [susuki:mezic:raak:hikihara:2016], quantum systems [klus:nuske:peitz:2022], robotics [mamakoukas:castano:tan:murphey:2019, bruder:fu:gillespie:remy:vasudevan:2020, haggerty:banks:kamenar:cao:curtis:mazic:hawkes:2023, rosenfelder:bold:eschmann:eberhard:worthmann:ebel:2024], and human-centered applications such as grip force prediction for robotic rehabilitation [bazina:kamenar:fonoberova:mezic:2025]. In EDMD, snapshots of the underlying system are employed to approximate the Koopman operator. It has been shown in [KordaMezi18] that EDMD converges in the infinite-data and infinite-dictionary (infinitely many observable functions) corresponding to a complexity solvability index two [colbrook2022foundations]. Then, first finite-data results on the estimation error for ergodic sampling for deterministic systems were provided in [Mezi22] followed by [ZhanZuaz23] analyzing also the $L^{2}$ projection error onto a finite element dictionary, cf. also the works [NuskPeit23, PhilScha24] providing error bounds for stochastic and control systems.
@@
More recently, kernel EDMD approximation schemes, which allow a data-informed choice of the dictionary, were proposed, see, e.g., [WillRowl15, klus:nuske:hamzi:2020]. Several works have already addressed $L^{2}$-error bounds for kernel EDMD: [PhilScha24:kernel] provides these in terms of the Mercer eigenvalues of the chosen kernel in the case of ergodic sampling along a trajectory, while the bounds in [PhilSchaWort2025, OlguOsseRami2025], are derived for i.i.d. sampling. The work [KohnPhil24] shows first pointwise (i.e., $L^{\infty}$) error bounds for approximations using Wendland kernels. However, while kernel EDMD is successfully used for stochastic systems, pointwise error bounds are still missing – to the best of the authors’ knowledge.
@@
In this paper, we propose a readily applicable Monte Carlo framework for estimation of kernel matrices to compute approximations of the Koopman operator for stochastic systems. Then, we provide an in-depth error analysis of the scheme. To this end, we establish Koopman invariance of (fractional) Sobolev spaces based on novel sufficient conditions on a Radon-Nikodym density associated with the stochastic dynamical system. Then we split the approximation error into a deterministic and a probabilistic part, where the latter results from Monte Carlo sampling of the expected value. Combining these ingredients, we then present novel full approximation error bounds for kernel EDMD for stochastic systems using Matérn and Wendland kernels. Finally, we present numerical simulations for systems governed by SDEs as a proof of concept.
@@
Notation For a set $\Omega\subset\mathbb{R}^{n}$, we denote the Borel sigma algebra on $\Omega$ (w.r.t. the standard topology) by $\mathcal{B}_{\Omega}$. The Lebesgue measure of a Borel set $\Delta\in\mathcal{B}_{\Omega}$ is denoted by $|\Delta|$; $dx$ or $dy$ indicate integrals w.r.t. Lebesgue measure. $\chi_{A}$ denotes the characteristic/indicator function of a set $A\subset\mathbb{R}^{n}$. The Radon-Nikodym density of a measure $\mu$ w.r.t. the Lebesgue measure will be denoted by $\frac{d\mu}{dx}$. The identity operator on some space is denoted by $\mathcal{I}$, the identity matrix especially is denoted by $I$. For Banach spaces $\mathcal{S}$ and $\mathcal{T}$, $L(\mathcal{S},\mathcal{T})$ denotes the space of bounded linear operators from $\mathcal{S}$ to $\mathcal{T}$. We write $L^{p}(\Omega)$, $p\in[1,\infty]$, for the usual Lebesgue spaces with respect to Lebesgue measure and $C_{b}(\Omega)$ for the space of bounded and continuous functions on $\Omega$; occasionally, we just write $L^{p}$ or $C_{b}$ if the domain is clear from context. For $f:\Omega\to\mathbb{R}$, we denote $\|f\|_{\infty}:=\sup_{x\in\Omega}|f(x)|$. The $\alpha$-th weak derivative of a function $f:\Omega\to\mathbb{R}$, $\alpha\in\mathbb{N}^{n}$, is denoted by $D^{\alpha}f$. In the case of a function $f:\Omega^{L}\to\mathbb{R}$, $L\in\mathbb{N}$, taking multiple arguments, we write $D_{l}^{\alpha}f$ for the $\alpha$-th derivative of $f$ “w.r.t. the $l$-th variable”, $l\in\{1,\dots,L\}$. For $d\in\mathbb{N}$, we abbreviate $\mathbbm{1}=\{v\in\mathbb{R}^{d}:|v_{i}|=1\}$.
@@ 2. Koopman operator for stochastic systems: kernel EDMD and error analysis
Let $\Omega\subset\mathbb{R}^{n}$ be a bounded domain with Lipschitz boundary and $\mathcal{B}_{\Omega}$ be the Borel sigma algebra on $\Omega$. Moreover, let $X$ and $Y$ be $\Omega$-valued random variables. It is well known (see, e.g., [Klen2008]) that there exists a function $\rho:\Omega\times\mathcal{B}_{\Omega}\to[0,1]$ such that
  1. i)
@@
A function with the first two properties is called a Markov transition kernel or Markov probability kernel; the third property makes $\rho$ a regular version of the conditional distribution of $Y$ given $X$ (see, e.g., [Kall2001]). In what follows, we consider stochastic dynamical systems of the form
@@ 2.1. The Koopman operator for stochastic systems
In the next chapter, we elaborate necessary and sufficient conditions for the invariance of Sobolev spaces $H^{s}(\Omega)$ under the Koopman operator. Those are of special interest since they arise as native spaces, or reproducing kernel Hilbert spaces, of certain kernels, for which there exists an extensive approximation framework, see Section 2.2, that allows for error bounds (Section 4). Before analyzing the Koopman operator on Sobolev spaces, we require the invariance of the space of square-integrable functions $L^{2}(\Omega)=H^{0}(\Omega)$ under $\mathcal{K}$.
@@
$\exists C>0:\quad\int\rho_{x}(\Delta)\,dx\,\leq\,C|\Delta|\qquad\,\forall\Delta\in\mathcal{B}_{\Omega},$ ($L^{1}$)
@@
Note that, by the Fubini-Tonelli theorem, ($L^{1}$) implies $L^{1}$-invariance of the Koopman operator. More generally, it can be shown that for a bounded domain $\Omega$, $L^{1}$-invariance (w.r.t. Lebesgue measure) implies $L^{q}$-invariance for $q>1$.
@@
While we consider $L^{2}(\Omega)=L^{2}_{{\boldsymbol{\lambda}}}(\Omega)$ with the Lebesgue measure $\boldsymbol{\lambda}$, our results can be extended to other reference measures – depending on the underlying stochastic system. In the case of $Y=X_{T}$ and $X=X_{0}$ for some Markov process $(X_{t})_{t\geq 0}$, the invariant measure $\mu$ of the process would be an intuitive choice, see, e.g., [NuskPeit23].
@@ 2.2. Recap on deterministic systems: kernel EDMD and error bound
with $F:\Omega\to\Omega$, which may be considered as a particular case of the stochastic case (1). In particular, we may consider the SDE (2) with $\sigma\equiv 0$ giving rise to the ordinary differential equation $\dot{x}=b(x)$. In the deterministic case, the transition kernel is given by
$\rho_{x}(A)=\delta_{F(x)}(A)$
such that the Koopman identity (3) reduces to
@@
i.e., $\mathcal{K}\psi=\psi\circ F$. In this case, it is often assumed that $F:\Omega\to\Omega$ is a diffeomorphism with $\inf_{x\in\Omega}|\det F^{\prime}(x)|=c>0$ and $\sup_{x\in\Omega}\|F^{\prime}(x)\|<\infty$ [KohnPhil24]. While the latter condition corresponds to the Lipschitz continuity of the drift function in (2), the assumption on the Jacobi determinant is a special case of ($L^{1}$). Specifically, applying the change of variables formula gives
@@
i.e., ($L^{1}$) also holds in the deterministic case.
We briefly recall Extended Dynamic Mode Decomposition (EDMD; [WillKevr15]) and its extension kernel EDMD, which are well-established numerical methods to approximate the Koopman operator $\mathcal{K}$ from data, see also [KutzBrun16] or the recent overview article [Colb24]. EDMD requires two ingredients: First, let $d\in\mathbb{N}$ snapshots $(x_{i},y_{i})$ with $y_{i}=F(x_{i})$, $i\in\{1,\dots,d\}$, be given. Then, we partition the snapshot data into the sets
@@
Second, let $M\in\mathbb{N}$ observable $L^{2}$-functions $\psi_{1},\dots,\psi_{M}$ be given, for which we use the notation $\Psi(x)=(\psi_{1}(x),\dots,\psi_{M}(x))^{\top}$ and define the data matrices
@@
The subspace $\mathcal{V}=\mathrm{span}\{\psi_{1},\dots,\psi_{M}\}\subset L^{2}(\Omega)$ spanned by the observables is a finite-dimensional space. From the Koopman identity, we get
@@
that is, the Koopman operator maps the components of $\Psi_{\mathcal{X}}$ to the components of $\Psi_{\mathcal{Y}}$. Hence, to derive a data-driven approximation of the Koopman operator, restricted to $\mathcal{V}$, we seek for an approximation $K\in\mathbb{R}^{M\times M}$ that satisfies this identity approximately, that is,
@@
Assuming that $\Psi_{\mathcal{X}}$ has full row rank (which is equivalent to strong linear independence of the observables $\psi_{1},\dots,\psi_{M}$ w.r.t. Lebesgue measure, see [PhilScha24, Lemma C.3]), an approximation can be computed as a solution to the least squares problem
@@
As illustrated above, EDMD crucially depends on two ingredients: a finite number of data pairs $\{(x_{i},y_{i})\mid y_{i}=F(x_{i}),\,i\in\{1,\dots,d\}\}$ and the choice of the observables. Convergence in the infinite-data limit has been shown in [KordaMezi18] and further specified w.r.t. the estimation error (number of data pairs $d$) in [Mezi22, NuskPeit23, PhilScha24]. The projection error is typically analyzed using finite-element observables as proposed in [ZhanZuaz23] and extended to control systems in [SchaWort23]. However, the latter only yields $L^{2}$-error bounds.
A particular variant of EDMD using a data-informed set of observables is kernel EDMD (kEDMD; [WillRowl15]), sharing two crucial features: First, from an application point of view, the choice of the observable functions is performed in a data-driven manner, eliminating fiddly and time-consuming tuning. Second, due to the rich mathematical structure of reproducing kernel Hilbert spaces, allows for pointwise deterministic error bounds.
@@
with the inner product $\langle\cdot,\cdot\rangle_{\mathbb{H}}$ holds. Then, the function $\Phi_{x}:\mathbb{R}^{n}\to\mathbb{R}$, $\Phi_{x}(y)=k(x,y)\in\mathbb{H}$, commonly termed the canonical feature of $k$ at $x\in\mathbb{R}^{n}$, plays a central role: We choose or sample points $\mathcal{X}=\{x_{1},\dots,x_{d}\}$ as before and let $y_{i}=F(x_{i})$. The dictionary is then given by the canonical features $\Phi_{x}=k(x,\cdot)$, $x\in\Omega$, of the data points, i.e.
$\mathcal{V}=\mathcal{V}_{\mathcal{X}}:=\mathrm{span}(\mathcal{D})=\mathrm{span}(\{\Phi_{x_{1}},\dots,\Phi_{x_{d}}\}).$
We immediately see that $M=d$ holds, i.e., we have exactly as many observables $\Phi_{x_{i}}$ as data points $x_{i}$. The entries in $\Psi_{X}\in\mathbb{R}^{d\times d}$ are easily calculated as $(\Psi_{X})_{i,j}=\psi_{i}(x_{j})=\Phi_{x_{i}}(x_{j})=k(x_{i},x_{j})$. The entries in $\Psi_{Y}$ are obtained via $(\Psi_{Y})_{i,j}=\psi_{i}(y_{j})=\Phi_{x_{i}}(F(x_{j}))=k(x_{i},F(x_{j}))$. Furthermore, we set
$\mathbf{K}_{\mathcal{X}}:=\mathbf{K}_{\mathcal{X},\mathcal{X}}=(k(x_{i},x_{j}))_{i,j=1}^{d}=\Psi_{X},\quad\mathbf{K}_{\mathcal{X},F(\mathcal{X})}=(k(x_{i},F(x_{j}))_{i,j=1}^{d}=\Psi_{Y}$
to which we will refer to as kernel matrices. Plugging these canonical features into (6) yields the regression problem
$\min_{K\in\mathbb{R}^{d\times d}}\left\|K\mathbf{K}_{\mathcal{X}}-\mathbf{K}_{\mathcal{X},F(\mathcal{X})}\right\|_{F}^{2}$ (9)
Assuming that the data points are pairwise distinct, $\mathbf{K}_{\mathcal{X}}$ is symmetric positive definite (hence invertible) due to the symmetry and strict positive definiteness of the underlying kernel $\mathbf{k}$. [^2] We explicitly speak of strict positive definiteness above, since positive definiteness of kernels is often defined via positive semi-definiteness of kernel matrices in the literature, see [KohnPhil24]. The regression problem (9) now reduces to (exactly) solving a linear system, whose solution $K=\mathbf{K}_{\mathcal{X},F(\mathcal{X})}\mathbf{K}_{\mathcal{X}}^{-1}$ corresponds to (7). This matrix not only represents the action of the Koopman operator for our data, but gives the matrix representation of $S_{\mathcal{X}}\mathcal{K}|_{\mathcal{V}}$ in the basis of the canonical features, where $S_{\mathcal{X}}:=S_{V_{\mathcal{X}}}$ denotes the orthogonal projection onto $\mathcal{V}$. Specifically, for some $g\in\mathbb{H}$, the projection onto $\mathcal{V}_{\mathcal{X}}$ is given by
@@
with $g_{\mathcal{X}}=(g(x_{1}),\dots,g(x_{d}))^{\top}$ and $\mathbf{k}_{\mathcal{X}}=(\Phi_{x_{1}},\dots,\Phi_{x_{d}})^{\top}$, see [KohnPhil24, Proposition 2.1]. For $f\in\mathcal{V}_{\mathcal{X}}$, we have some $\alpha_{f}\in\mathbb{R}^{d}$ such that $f=\alpha_{f}^{\top}\mathbf{k}_{\mathcal{X}}$. Setting $g=\mathcal{K}|_{\mathcal{V}_{\mathcal{X}}}f$, we get $g_{\mathcal{X}}=\mathbf{K}_{\mathcal{X},F(\mathcal{X})}^{\top}\alpha_{f}$ and
@@
This means the kEDMD-approximant $K=\mathbf{K}_{\mathcal{X},F(\mathcal{X})}\mathbf{K}_{\mathcal{X}}^{-1}$ [KohnPhil24, Proposition 3.2] maps (the coefficients of) $f\in\mathcal{V}_{\mathcal{X}}$ to (the coefficients of) $S_{\mathcal{X}}\mathcal{K}|_{\mathcal{V}_{\mathcal{X}}}f$ linearly (in the respective canonical bases). To extend this operator to the entire space $\mathbb{H}$, one may simply add another orthogonal projection from the right, i.e., replace $f\in\mathcal{V}_{\mathcal{X}}$ by $S_{\mathcal{X}}g$ for any $g\in\mathbb{H}$. From (10), we get the representation $S_{\mathcal{X}}g=\alpha_{S_{\mathcal{X}}g}^{\top}\mathbf{k}_{\mathcal{X}}=g_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}$. In total, this yields the Koopman approximant
@@
The main result regarding this approximation obtained in [KohnPhil24, Theorem 3.4] may be summarized as follows.
@@
    Invariance property. First, it is assumed that the Koopman operator is indeed a bounded operator on $\mathbb{H}$. Due to closedness of the Koopman operator on RKHS, it suffices to show that $\mathcal{K}\mathbb{H}\subset\mathbb{H}$ (see [KohnPhil24, Lemma 3.1] and Proposition 2.4 below).

### v2

@@ 1. Introduction
Nowadays, the Koopman operator is an established mathematical and computational tool for modeling, analysis, and control of dynamical systems, see, e.g., the collection [30] or the recent overview papers [10, 47]. The underlying idea is to consider nonlinear dynamics through the lens of observable functions using the linear, but infinite-dimensional Koopman operator, see the original work by B.O. Koopman [26, 25] and the seminal paper [32] by I. Mezić. This enabled, among others, to leverage modern techniques from spectral analysis to investigate the complex dynamical behaviour of the original nonlinear system [44].
@@
Recently, data-driven approaches like dynamic mode decomposition (EDMD; [53, 11]) were proposed to approximate the Koopman operator, see, e.g., the recent overview [13], which has lead to a plethora of successful applications, including fluid flows [4, 33], climate forecasting [17, 3], molecular dynamics [35], power systems technology [48], quantum systems [23], robotics [29, 9, 16, 43], and human-centered applications such as grip force prediction for robotic rehabilitation [5]. In EDMD, snapshots of the underlying system are employed to approximate the Koopman operator. It has been shown in [27] that EDMD converges in the infinite-data and infinite-dictionary (infinitely many observable functions) limit, corresponding to a complexity solvability index of two [12]. Then, first finite-data results on the estimation error for ergodic sampling for deterministic systems were provided in [34] followed by [54] analyzing the $L^{2}$-projection error onto a finite element dictionary, cf. also the works [36, 39] providing error bounds for stochastic and control systems.
@@
More recently, kernel EDMD approximation schemes, which allow a data-informed choice of the dictionary, were proposed, see, e.g., [52, 22]. Several works have already addressed $L^{2}$-error bounds for kernel EDMD: [40] provides these in terms of the Mercer eigenvalues of the chosen kernel in the case of ergodic sampling along a trajectory, while the bounds in [41, 38], are derived for i.i.d. sampling. The work [24] shows first pointwise (i.e., $L^{\infty}$) error bounds for approximations using Wendland kernels. However, while kernel EDMD is successfully used for stochastic systems, pointwise error bounds are still missing – to the best of the authors’ knowledge.
@@
In this paper, we propose a readily applicable Monte Carlo framework for estimation of kernel matrices to compute approximations of the Koopman operator for stochastic systems. Then, we provide an in-depth error analysis of the scheme. To this end, we establish Koopman invariance of (fractional) Sobolev spaces based on novel sufficient conditions on a Radon-Nikodym density associated with the stochastic dynamical system. Then we split the approximation error into a deterministic and a probabilistic part, where the latter results from Monte Carlo estimation of the expected value. Combining these ingredients, we then present novel full approximation error bounds for kernel EDMD for stochastic systems using Matérn and Wendland kernels. Finally, we present numerical simulations for systems governed by SDEs as a proof of concept.
@@
Notation. For a set $\Omega\subset\mathbb{R}^{n}$, we denote the Borel sigma algebra on $\Omega$ (w.r.t. the standard topology) by $\mathcal{B}_{\Omega}$. The Lebesgue measure of a Borel set $\Delta\in\mathcal{B}_{\Omega}$ is denoted by $|\Delta|$; $dx$ or $dy$ indicate integrals w.r.t. Lebesgue measure. $\chi_{A}$ denotes the characteristic/indicator function of a set $A\subset\mathbb{R}^{n}$. The Radon-Nikodym density of a measure $\mu$ w.r.t. the Lebesgue measure will be denoted by $\frac{d\mu}{dx}$. The identity operator on some space is denoted by $\mathcal{I}$, the identity matrix especially is denoted by $I$. For Banach spaces $\mathcal{X}$ and $\mathcal{Y}$, $L(\mathcal{X},\mathcal{Y})$ denotes the space of bounded linear operators from $\mathcal{X}$ to $\mathcal{Y}$. The norm of $T\in L(\mathcal{X},\mathcal{Y})$ is denoted by $\|T\|_{\mathcal{X}\to\mathcal{Y}}$. We write $L^{p}(\Omega)$, $p\in[1,\infty]$, for the usual Lebesgue spaces with respect to Lebesgue measure, $C_{b}(\Omega)$ for the space of bounded and continuous functions, and $C_{c}^{\infty}(\Omega)$ for the space of infinitely differentiable and compactly supported functions, respectively, on $\Omega$; occasionally, we just write $L^{p}$, $C_{b}$, or $C_{c}^{\infty}$ if the domain is clear from context. For $f:\Omega\to\mathbb{R}$, we denote $\|f\|_{\infty}:=\sup_{x\in\Omega}|f(x)|$. The $\alpha$-th weak derivative of a function $f:\Omega\to\mathbb{R}$, $\alpha\in\mathbb{N}^{n}$, is denoted by $D^{\alpha}f$. In the case of a function $f:\Omega^{L}\to\mathbb{R}$, $L\in\mathbb{N}$, taking multiple arguments, we write $D_{l}^{\alpha}f$ for the $\alpha$-th derivative of $f$ “w.r.t. the $l$-th variable”, $l\in\{1,\dots,L\}$.
@@ 2. Koopman operator for stochastic systems: kernel EDMD and error analysis
Let $\Omega\subset\mathbb{R}^{n}$ be a bounded domain with Lipschitz boundary and $\mathcal{B}_{\Omega}$ be the Borel sigma algebra on $\Omega$. Moreover, let $X$ and $Y$ be $\Omega$-valued random variables. It is well known (see, e.g., [20]) that there exists a function $\rho:\Omega\times\mathcal{B}_{\Omega}\to[0,1]$ such that
  1. (i)
@@
A function with the first two properties is called a (Markov) transition kernel or probability kernel; the third property makes $\rho$ a regular version of the conditional distribution of $Y$ given $X$ (see, e.g., [19]). In what follows, we consider stochastic dynamical systems of the form
@@ 2.1. The Koopman operator for stochastic systems
In the next section, we elaborate necessary and sufficient conditions for the invariance of Sobolev spaces $H^{s}(\Omega)$ under the Koopman operator. These spaces are of special interest since they arise as native spaces, or reproducing kernel Hilbert spaces (RKHS), of certain kernels for which there exists an extensive approximation framework, see Section 2.2, that allows for error bounds (Section 4). Before analyzing the Koopman operator on Sobolev spaces, we require the invariance of the space of square-integrable functions $L^{2}(\Omega)=H^{0}(\Omega)$ under $\mathcal{K}$.
@@
$\exists C>0:\quad\int\rho_{x}(\Delta)\,dx\,\leq\,C|\Delta|\qquad\,\forall\Delta\in\mathcal{B}_{\Omega},$ (4)
@@
Note that, by the Fubini-Tonelli theorem, (4) implies $L^{1}$-invariance of the Koopman operator. More generally, it can be shown that for a bounded domain $\Omega$, $L^{1}$-invariance (w.r.t. Lebesgue measure) implies $L^{q}$-invariance for $q>1$.
@@
While we consider $L^{2}(\Omega)=L^{2}_{{\boldsymbol{\lambda}}}(\Omega)$ with the Lebesgue measure $\boldsymbol{\lambda}$, our results can be extended to other reference measures – depending on the underlying stochastic system. In the case of $Y=X_{T}$ and $X=X_{0}$ for some Markov process $(X_{t})_{t\geq 0}$, the invariant measure $\mu$ of the process would be an intuitive choice, see, e.g., [36].
@@ 2.2. Recap on deterministic systems: kernel EDMD and error bound
with $F:\Omega\to\Omega$, which may be considered as a particular case of the stochastic case (1). In particular, we may consider the SDE (2) with $\sigma\equiv 0$ giving rise to the ordinary differential equation $\dot{x}=b(x)$. In this case, the transition kernel is given by $\rho_{x}(A)=\delta_{F(x)}(A)$, such that the Koopman identity (3) reduces to
@@
i.e., $\mathcal{K}\psi=\psi\circ F$. As in [24], we assume that $F:\Omega\to\Omega$ is a diffeomorphism satisfying $\inf_{x\in\Omega}|\det F^{\prime}(x)|=c>0$ and $\sup_{x\in\Omega}\|F^{\prime}(x)\|<\infty$ . While the latter condition corresponds to the Lipschitz continuity of the drift function in (2), the assumption on the Jacobi determinant is a special case of (4). Applying the change of variables formula gives
@@
i.e., (4) also holds in the deterministic case.
We briefly recall Extended Dynamic Mode Decomposition (EDMD; [53]) and its extension kernel EDMD, which are well-established numerical methods to approximate the Koopman operator $\mathcal{K}$ from data, see also [28] or the recent overview article [13]. EDMD requires two ingredients: First, let $d\in\mathbb{N}$ snapshots $(x_{i},y_{i})$ with $y_{i}=F(x_{i})$, $i\in\{1,\dots,d\}$, be given. Then, we partition the snapshot data into the sets
@@
Second, let $M\in\mathbb{N}$ observable $L^{2}$-functions $\psi_{1},\dots,\psi_{M}$ be given, for which we use the notation $\Psi(x)=(\psi_{1}(x),\dots,\psi_{M}(x))^{\top}$, and define the data matrices
@@
The subspace $\mathcal{V}=\mathrm{span}\{\psi_{1},\dots,\psi_{M}\}\subset L^{2}(\Omega)$ spanned by the observables is finite-dimensional. From the Koopman identity, we get
@@
that is, the Koopman operator maps the components of $\Psi_{\mathcal{X}}$ to the components of $\Psi_{\mathcal{Y}}$. Since the Koopman operator acts linearly on observables, we seek a data-driven linear approximation on $\mathcal{V}$ given by a matrix $K\in\mathbb{R}^{M\times M}$ approximately satisfying (5), that is,
@@
Assuming that $\Psi_{\mathcal{X}}$ has full row rank (which is equivalent to strong linear independence of the observables $\psi_{1},\dots,\psi_{M}$ w.r.t. Lebesgue measure, see [39, Lemma C.3]), an approximation can be computed as a solution to the least squares problem
@@
As illustrated above, EDMD crucially depends on two ingredients: a finite number of data pairs $\{(x_{i},y_{i})\mid y_{i}=F(x_{i}),\,i\in\{1,\dots,d\}\}$ and the choice of the observables. Convergence in the infinite-data limit has been shown in [27] and further specified w.r.t. the estimation error (number of data pairs $d$) in [34, 36, 39]. The projection error is typically analyzed using finite-element observables as proposed in [54] and extended to control systems in [46]. However, the latter only yields $L^{2}$-error bounds.
A particular variant of EDMD using a data-informed set of observables is kernel EDMD (kEDMD; [52]), sharing two crucial features: First, from an application point of view, the choice of the observable functions is performed in a data-driven manner, eliminating fiddly and time-consuming tuning. Second, the rich mathematical structure of reproducing kernel Hilbert spaces allows for pointwise deterministic error bounds.
@@
holds, where $\Phi_{x}\in\mathbb{H}$, defined by $\Phi_{x}(y)=\text{\sf{k}}(x,y)$, is commonly termed the canonical feature of k at $x\in\Omega$. In kEDMD, these features play a central role: We choose or sample points $\mathcal{X}=\{x_{1},\dots,x_{d}\}$ as before and let $y_{i}=F(x_{i})$. The dictionary is given by the canonical features $\Phi_{x}=\text{\sf{k}}(x,\cdot)$, $x\in\Omega$, at the data points, i.e.
$\mathcal{V}=\mathcal{V}_{\mathcal{X}}:=\mathrm{span}(\mathcal{D})=\mathrm{span}\{\Phi_{x_{1}},\dots,\Phi_{x_{d}}\}.$
Clearly $M=d$ holds, i.e., we have exactly as many observables $\Phi_{x_{i}}$ as data points $x_{i}$. The data matrix entries are easily calculated as $(\Psi_{\mathcal{X}})_{i,j}=\psi_{i}(x_{j})=\Phi_{x_{i}}(x_{j})=\text{\sf{k}}(x_{i},x_{j})$ and $(\Psi_{\mathcal{Y}})_{i,j}=\psi_{i}(y_{j})=\Phi_{x_{i}}(F(x_{j}))=\text{\sf{k}}(x_{i},F(x_{j}))$. We write them as kernel matrices
$\mathbf{K}_{\mathcal{X}}:=\mathbf{K}_{\mathcal{X},\mathcal{X}}=(\text{\sf{k}}(x_{i},x_{j}))_{i,j=1}^{d}=\Psi_{\mathcal{X}},\qquad\mathbf{K}_{\mathcal{X},F(\mathcal{X})}:=(\text{\sf{k}}(x_{i},F(x_{j}))_{i,j=1}^{d}=\Psi_{\mathcal{Y}}.$
Plugging the kernel matrices into (7) yields the regression problem
$\min_{K\in\mathbb{R}^{d\times d}}\left\|K\mathbf{K}_{\mathcal{X}}-\mathbf{K}_{\mathcal{X},F(\mathcal{X})}\right\|_{F}^{2}.$ (10)
Assuming that the data points are pairwise distinct, $\mathbf{K}_{\mathcal{X}}$ is symmetric positive definite (hence invertible) due to the symmetry and strict positive definiteness of the underlying kernel $\mathbf{k}$. [^2] We explicitly speak of strict positive definiteness above, since positive definiteness of kernels is often defined via positive semi-definiteness of kernel matrices in the literature, see [24]. The regression problem (10) now reduces to (exactly) solving a linear system, whose solution $K=\mathbf{K}_{\mathcal{X},F(\mathcal{X})}\mathbf{K}_{\mathcal{X}}^{-1}$ corresponds to (8). This matrix not only represents the action of the Koopman operator for our data, but gives the matrix representation of $S_{\mathcal{X}}\mathcal{K}|_{\mathcal{V}}$ in the basis of the canonical features, where $S_{\mathcal{X}}:=S_{V_{\mathcal{X}}}$ denotes the orthogonal projection onto $\mathcal{V}$. Specifically, for some $g\in\mathbb{H}$, the projection onto $\mathcal{V}_{\mathcal{X}}$ is given by
@@
with $g_{\mathcal{X}}:=(g(x_{1}),\dots,g(x_{d}))^{\top}$ and $\mathbf{k}_{\mathcal{X}}:=(\Phi_{x_{1}},\dots,\Phi_{x_{d}})^{\top}$, see [24, Proposition 2.1]. For $f\in\mathcal{V}_{\mathcal{X}}$, there exists $\alpha_{f}\in\mathbb{R}^{d}$ such that $f=\alpha_{f}^{\top}\mathbf{k}_{\mathcal{X}}$. With $g=\mathcal{K}|_{\mathcal{V}_{\mathcal{X}}}f$ we get $g_{\mathcal{X}}=\mathbf{K}_{\mathcal{X},F(\mathcal{X})}^{\top}\alpha_{f}$ and
@@
This means the kEDMD-approximant $K=\mathbf{K}_{\mathcal{X},F(\mathcal{X})}\mathbf{K}_{\mathcal{X}}^{-1}$ [24, Proposition 3.2] maps (the coefficients of) $f\in\mathcal{V}_{\mathcal{X}}$ to (the coefficients of) $S_{\mathcal{X}}\mathcal{K}|_{\mathcal{V}_{\mathcal{X}}}f$ linearly (in the respective canonical bases). To extend this operator to the entire space $\mathbb{H}$, one may simply add another orthogonal projection from the right, i.e., replace $f\in\mathcal{V}_{\mathcal{X}}$ by $S_{\mathcal{X}}g$ for any $g\in\mathbb{H}$. From (11), we get the representation $S_{\mathcal{X}}g=\alpha_{S_{\mathcal{X}}g}^{\top}\mathbf{k}_{\mathcal{X}}=g_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}$. In total, this yields the Koopman approximant
@@
The main result regarding this deterministic Koopman approximation [24, Theorem 3.4] may be summarized as follows.
@@
    Invariance property. First, it is assumed that the Koopman operator is indeed a bounded operator on $\mathbb{H}$. Due to closedness of the Koopman operator on RKHS, it suffices to show that $\mathcal{K}\mathbb{H}\subset\mathbb{H}$ (see [24, Lemma 3.1] and Proposition 2.4 below).